2.5 is 40% of what number?
Found 2 solutions by stanbon, london maths tutor:
Answer by stanbon(75887)

2.5 is 40% of what number?
---
Let the number be "x":
Equation:
2.5 = 0.40x
x = 2.5/0.40
x = 6.25
================
Cheers,
Stan H.

Answer by london maths tutor(243)

2.5 is 40% of what number?
40% = 40/100
Let's the number be "x"
(40/100)*x = 2.5
40x= 2.5*100
40x = 250
x = 250/40
x = 25/4
x = 6 1/4 or x = 6.25 (answer as decimal)